OpenStudy (dtan5457):

What is the volume occupied by 3.00 moles of nitrogen gas (n2) at STP?

OpenStudy (dtan5457):

So I believe you have to do the ideal gas law. p=1atm v=??? t=273k r=0.0821 (molar gas constant) n=3 so in this case..I would think it would be v x 3=273(0.0821)(1atm) I get v=7.47 liters but my book says its 67.2 what did i do wrong?

OpenStudy (dtan5457):

@aaronq

OpenStudy (dtan5457):

@Kainui @Abhisar

OpenStudy (anonymous):

You simply use the rule aside from Avogadros Contant, in which 1mol of a substance = 22.4litres . What your equation is is 3mol=x. Then x=3*22.4/1=67.2litres. Got it?

OpenStudy (dtan5457):

wow never mind....forgot to do some reading on my book. standard mole has 22.4 L as the volume...i just multiply by 3 here.